### v2

# results on dynamics of bungee set of composite entire functions in the eremenko-lyubich class

Abstract.
In this paper, we have discussed the dynamics of composite entire functions in terms of relationship between bungee set, repelling periodic points (to be denoted by $RP$) and rationally indifferent fixed point set. We have established relation between the dynamics of composite function and the functions taken for composition. We showed that the union of their Bungee sets contains Bungee set of the composite function. We also showed that $RP$ set of composite functions contains the $RP$ set of functions used for the composition. We have mostly dealt with the permutable functions of class $\mathcal{B}$. These results hold true when we have a composite function in which one function is a periodic translation of another.
Key words and phrases:
bungee set, wandering domain, Baker domain, repelling periodic point, rationally indifferent fixed point, escaping set, asymptotic value, singular set
2020 Mathematics Subject Classification:
37F10, 30D05
The research work of the first author is supported by ANRF(SERB) research grant TAR/2023/000197
1. Introduction
When we iterate a homomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n-$times, we observe that the points of the complex plane move differently, which we call the dynamics of the point where we denote $f^{n}$ to be $n-$th iteration of a function, where $n\in\mathbb{N}$. As, not all points have same dynamics, we therefore have different classifications of points of the complex plane $\mathbb{C}$; one classification is based on the nature of neighbourhood of a point $z_{0}$ (say); they are said to be in the Fatou set $F(f)$ (set of points whose neighbourhood exhibits stable behaviour) and Julia set $J(f)$ (set of points whose neighbourhood exhibits chaotic behaviour). This was broadly the classification we generally see when we talk about dynamics of points in complex plane.
There is yet another classification of these points based on the nature of their orbits; viz., the Filled in Julia set $K(f)$ (set of points whose orbits are bounded), escaping set $I(f)$ (set of points whose orbits tend to infinity), and bungee set $BU(f)$ (the orbits of such points neither remain bounded nor tend to infinity; only there exists a subsequence that remains bounded and another one that is unbounded). These sets are disjoint among themselves and their union $BU(f)\cup I(f)\cup K(f)=\mathbb{C}$. In [16], it was shown that if a Fatou component intersects with $BU(f)$ then that component becomes a wandering domain contained in $BU(f)$. $K(f)$ comprises of Fatou components in which orbits under iteration stays bounded, along with repelling periodic points and rationally indifferent periodic points that lies inside Julia set $J(f).$ It is to be noted that the bungee set and escaping set must have a finite intersection with $J(f)$ (see [5, 16]). The escaping set too has a finite intersection with the Fatou set when the Fatou component is a Baker domain. Knowing all these sets for an entire function, we were curious about what happens when we compose two functions and how these sets relate to the composition with those of individual functions. We tried to see what happens to the dynamics of transcendental functions under compositions as composition of two transcendental functions is also transcendental.
Simplifying the situation further, we have taken transcendental functions $f,g$ of Eremenko Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded). Additionally, this class does not have any intersection of the Fatou set with the bungee set, in other words there is no wandering domain which also lies inside the bungee set for $f\in\mathcal{B}.$ Since our aim was to establish a relationship between different sets of functions and their composition and it would be quite difficult to establish any relation where there is no algebraic relation among these functions, we have focused on commuting(permutable) functions, that is, $f\circ g=g\circ f$, and $g=f+C$ where $C$ is the period of $f$.
The following results are proven in this paper:
Let $f,g\in\mathcal{B}$ be commuting functions. The bungee set of composed functions is contained in the union of the bungee set of functions used for composition. The set of repelling periodic and indifferent fixed points of the composed function $RP(f\circ g)$ contains the union of repelling periodic and indifferent fixed points of these functions; that is, $RP(f\circ g)\supset RP(f)\cup RP(g)$.
We know some results for exponential functions which belong to Speiser class $\mathcal{S}$ (such class of functions has finite number of elements in their singular set) [18]. Also, $\mathcal{S}\subset\mathcal{B}$. $f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes under composition arewith $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$ [2]; therefore, for such a case, all sets of composed functions are equal to the functions taken for composition.
1.1. Observations
While developing the relationships among the functions and the individual functions, we also found that the bungee set and escaping set exhibit similar characteristics for functions of class $\mathcal{B}$:
  1. (1)
    The closure of both these sets equals Julia set.
  2. (2)
    Both sets are neither open nor closed (3.2).
  3. (3)
    Both sets of composite function is contained in union of individual function (to be proved in this paper).
2. Main results
To begin with the commuting entire functions, we first understand the nature of these sets when functions have no asymptotic values in their singular set.
Theorem 2.1.
If $f$ and $g$ are commuting functions having no finite asymptotic values, then $I(f)$ is completely invariant under $g$ and vice versa.
Proof.
[…]
But assuming $g(z)$ does not have any finite asymptotic value, so $|g(f^{n}(z_{0})|\rightarrow\infty$.
Since $f\,\&\,g$ are commuting so, $|g(f^{n}(z_{0})|$ = $|f^{n}(g(z_{0})|\rightarrow\infty.$ This implies that $g(z)$ carries $I(f)$ inside itself or in other words, $I(f)$ is forward invariant under $g$. It is easy to see that $g^{-1}(I(f))\subset I(f)$. Hence, $I(f)$ is completely invariant under $g$. As $f\,\&\,g$ are commuting, their roles can be interchanged and hence, $I(g)$ is completely invariant under $f$ as well. Hence, $I(f)$ and $I(g)$ is completely invariant under $f\circ g$ as well. ∎
Theorem 2.2.
Suppose $f$ and $g$ are commuting entire functions having no asymptotic values. Then $K(f)$ and $BU(f)$ are completely invariant under $g$.
Proof.
$g(K(f))$ is, in general forward invariant, because if we assume a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|f^{m_{k}}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$, thereby proving the result. Now, let $z_{0}\in g^{-1}(K(f)).$ This implis that $g(z_{0})\in K(f)$ and $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ As these functions do not have any finite asymptotic values, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$. Thus, we have that $K(f)$ is completely invariant under $g$ when the functions have no finite asymptotic values.
Now bungee set, $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know that $I(f)$ and $K(f)$ are completely invariant under $g$ (2.1) and hence, $BU(f)$ is also completely invariant under $g$. ∎
Lemma 2.3.
The filled-in Julia set of the composite function $f\circ g$ cannot be larger than intersection of filled in Julia set of the individual functions having no finite asymptotic values.
Proof.
We assume that $K(f\circ g)\supset K(f)\cap K(g)$. Thus, there might be points in $K(f\circ g)$ that have a finite intersection with $I(f)\cup I(g)\cup BU(f)\cup BU(g)\cup A\cup B$ where $A:K(f)\backslash K(f)\cap K(g)$ and $B:K(g)\backslash K(f)\cap K(g)$. If $z_{0}\in I(f)\cup I(g)\cup BU(f)\cup BU(g)$, these points cannot be bounded as one function carries the points of other functions to itself, meaning $f^{n}\circ g^{n}(z_{0})$ as $n\rightarrow\infty$ will either remain oscillatory like bungee points nor will tend to $\infty$. Now, let $w_{0}\in A$ and apply $f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ as $m\rightarrow\infty$, which is bounded because $g(K(f))\subset K(f)$. However, $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$, which cannot be bounded ($f(BU(g)\subset BU(g)$ and $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ cannot be bounded. As these functions have no finite asymptotic value, so $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ will be unbounded in the case when $w_{0}\in I(g).$) Similarly, such arguments may be given when $w_{1}\in B$. Thus, we arrive at a contradiction. Hence, $K(f\circ g)\subset K(f)\cap K(g)$. ∎
Theorem 2.4.
Commuting transcendental entire functions with no finite asymptotic values have equal bungee sets, filled-in Julia sets and escaping sets.
Proof.
Let $z_{0}\in BU(f\circ g)$ where $f\ \&\ g$ do not have any finite asymptotic value. To begin with let there be two subsequences $\{m_{k}\}$ and $\{n_{k}\}$ both tending to $\infty$ such that $|(f\circ g)^{m_{k}}(z_{0})|<R$ for some $R$ and $|(f\circ g)^{n_{k}}(z_{0})|\rightarrow\infty.$
So, $|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R$ for some $R.$ As $f$ does not have any finite asymptotic value, so $|g^{m_{k}}(z_{0})|<R_{1}$ for some $R_{1}$. Similarly for $f,$ since they are commuting so, $|f^{m_{k}}(z_{0})|<R_{2}$ for some $R_{2}$.
Now, $|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty.$ We have 3 possible cases: $|g^{n_{k}}(z_{0})|\rightarrow\infty$ or $|g^{n_{k}}(z_{0})|\rightarrow R_{3}$ for some $R_{3}$ or $g^{n_{k}}(z_{0})$ exhibits bungee behaviour. If we have the second case, it means $z_{0}\in K(g)$ (This will imply all subsequences of ($g^{j}(z_{0})$) is bounded where $j$ is an arbitrary index). To understand this, we take $w_{0}\in K(g)$. We observe $g^{j}\circ f^{2j}\circ g^{j}(w_{0})\in g^{j}\circ f^{2j}(K(g))\subset g^{j}(K(g))$ which is bounded for $j\rightarrow\infty$. Hence, this case isn’t possible. Now, when we have the third and the first case, then both essentially states that $z_{0}\in BU(g)$. From (2.2), we have $g^{n_{1k}}\circ f^{2n_{1k}}\circ g^{n_{1k}}(z_{0})$ exhibits bungee behaviour which is as expected. And due to commutation, we also conclude that $z_{0}\in BU(f)$ as well. Hence, $BU(f\circ g)\subset BU(f)\cap BU(g)$.
Given the preceding proof, we can ask a question:
Question: What can we say about points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g),$ where will it be situated?
Answer: As the functions $f$ and $g$ do not have any finite asymptotic values, therefore, points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g)$ cannot intersect with $K(f\circ g).$ Also, $I(f\circ g)\subset I(f)\cup I(g)$ [14] and as $(BU(f)\cup BU(g))\cap(I(f)\cup I(g))=\emptyset,$ it means that $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g)$ cannot intersect with $I(f\circ g)$ either. This implies that the only place it will be contained is $BU(f\circ g)$. However, if it happens, then such points must be contained in $BU(f)\cap BU(g)$ implying that no such points exist. This is only possible if $BU(f)=BU(g)$.
Similar question can be asked for filled-in Julia sets as well, for, it was shown in Lemma 2.3 that $K(f\circ g)\subset K(f)\cap K(g)$. And the answer is: Let $w_{0}\in K(f\circ g)\ \backslash\ K(f)\cap K(g)$. These points won’t intersect with $I(f\circ g)$ and neither with $BU(f\circ g)$ as $BU(f\circ g)=BU(f)=BU(g)$ and $BU(f\circ g)\cap K(f)\cup K(g)=\emptyset.$ Similarly, for intersection of filled-in Julia sets with escaping sets. Hence, we conclude no such points are possible. Therefore, $K(f\circ g)=K(f)=K(g)$ when there are no finite asymptotic values for these commuting functions.
So given the following results:
$K(f\circ g)$ $=K(f)=K(g)$
and
(2.1) $BU(f\circ g)$ $=BU(f)=BU(g)$
one can conclude that $I(f\circ g)=I(f)=I(g)$, for we know that $K(f)\cup I(f)\cup BU(f)=\mathbb{C}=K(f\circ g)\cup I(f\circ g)\cup BU(f\circ g)=K(g)\cup I(g)\cup BU(g)$ and from the previous results (2) we can conclude that $I(f\circ g)=I(f)=I(g)$. ∎
The following examples illustrate the results established above.
Example 2.5.
Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ do not have any finite asymptotic values
One may also consider the following functions for illustration of above result:
$f_{1}(z)=1+z+e^{z}$ which would be commuting with $g_{1}(z)=1+z+e^{z}+2\pi i$; $f_{2}(z)=z+sin(z)$ which would be commuting with $g_{2}(z)=z+sin(z)+2\pi.$
3. Some results for functions of class $\mathcal{B}$
Now, we will prove some results when these functions indeed have some finite asymptotic values, and to simplify, we will focus on the functions of the Eremenko-Lyubich class $\mathcal{B}$. These functions do not have a Baker domain; therefore, the escaping set does not intersect with the Fatou set. In addition, they do not have a finite intersection of the Bungee set with the wandering domain because the wandering domain of such functions does not have (sub)sequences that tend to $\infty$.
Therefore, the Julia set for functions of class $\mathcal{B}$ comprises of $RP(f),BU(f)$ and $I(f)$; that is, $J(f)=BU(f)\cup I(f)\cup RP(f)$. For commuting functions $f$ and $g$, we have $J(f\circ g)\supset J(f)\cup J(g)$ [10, 11]. However, when these are commuting functions of class $\mathcal{B}$ we have, $J(f)=J(g)=J(f\circ g)$ [10]. We also know that $g(J(f))=J(f)$ and $\overline{RP(f)}=J(f),\overline{I(f)}=J(f)$ [4, 6].
Proposition 3.1.
The bungee set for functions of class $\mathcal{B}$ does not intersect with Fatou set.
Proof.
[…]
This also means that $BU(f)\subset J(f)$ only for functions of class $\mathcal{B}.$ Also, it is easy to see that $BU(f)$ is completely invariant by the complete invariance of $I(f)$ and $K(f)$. As a result, by the minimality property of Julia set, $\overline{BU(f)}=J(f)$.
The escaping set of a function is in general, neither open, nor closed [5]. We have similar kind of result for the bungee set.
Proposition 3.2.
The bungee set for functions of class $\mathcal{B}$ is neither open nor closed.
Proof.
We can easily prove the complete invariance of bungee set by noting that $BU(f)=\mathbb{C}\backslash(I(f)\cup K(f))$. Now, if we assume that bungee set is closed then it will be a closed completely invariant set. This implies that $J(f)\subset BU(f)$ (because $J(f)$ is the smallest closed completely invariant set) which is a contradiction. Furthermore, if bungee set is open, it will imply that it has interior to it which is also not possible (because $BU(f)\subset J(f)$ and having interior of bungee set means Julia set also has interior which is not possible for an entire function $f$ for which $F(f)\neq\emptyset$ [3]). Hence, bungee set is neither open nor closed. ∎
Theorem 3.3.
Functions of class $\mathcal{B}$ which are periodic translate of each other also have equal bungee sets, escaping sets and filled-in Julia sets.
Proof.
We take $f(z)$ and $g(z)=f(z)+C$ where $C$ is the periodic translate of $f$ (more generally, we can take $g=f^{l}(z)+C$, for some $l\in\mathbb{N}$). Then $g^{n}(z)=f^{n}(z)+C$ which is almost commuting. Here we can see that if a point is bounded under $f^{n},$ it is also bounded under $g^{n}.$ Similarly, if a point is escaping under $f^{n},$ it is also escaping under $g^{n}$. Hence, $BU(f)=BU(g),I(f)=I(g),K(f)=K(g)$. ∎
Such functions can be viewed as almost commuting type with a constant shift.
Now moving forward, we will study functions that are exactly commuting of class $\mathcal{B}$ and having finite asymptotic values.
Lemma 3.4.
$g(BU(f))\subset BU(f)\cup RP(f)$ for functions in class $\mathcal{B}.$
Proof.
If we recall the definition of a bungee point, there must exist a subsequence which is bounded and another subsequence which is unbounded. Now it is obvious that when the subsequence which is bounded $f^{m}(z_{0})\rightarrow l_{0}$ (say) and when acted upon by $g$ will also keep it bounded (since $f,g$ are entire functions). Let us assume that there is only one such bounded subsequence and $l$ unbounded subsequence which are as mentioned below.
$f^{m}(z_{0})\rightarrow l_{0}$
$f^{n_{1}}(z_{0})\rightarrow\infty$
$f^{n_{2}}(z_{0})\rightarrow\infty$
$\vdots$
$f^{n_{l}}(z_{0})\rightarrow\infty$
Now, there are two possible cases when $g$ acts on those unbounded subsequences.
Case 1. All these subsequence gets bounded, i.e.,
$g(f^{n_{1}}(z_{0}))\rightarrow w_{1}\text{ which also means }f^{n_{1}}(g(z_{0}))\rightarrow w_{1}$
$g(f^{n_{2}}(z_{0}))\rightarrow w_{2}\text{ which also means }f^{n_{2}}(g(z_{0}))\rightarrow w_{2}$
$\vdots$
$g(f^{n_{l}}(z_{0}))\rightarrow w_{l}\text{ which also means }f^{n_{l}}(g(z_{0}))\rightarrow{w_{l}}$
If this happens, then $g(z_{0})\in$ $K(f)=F(f)\cup RP(f)$. As $z_{0}\in J(f)$ as well and we know that $g(J(f))=J(f)$ hence, $g(z_{0})\in RP(f)$ for the above case. Now, let us understand what does this mean by $g(z_{0})\in RP(f)$: It means that $f^{p}(g(z_{0}))=g(z_{0})$ for some $p\in\mathbb{N}.$ This also means $g(f^{p}(z_{0}))=g(z_{0})$. Because $f$ and $g$ are not invertible functions we have $f^{p}(z_{0})\not=g(z_{0})$. So, $g(z_{0})\in RP(f)\ \&\ z_{0}\in BU(f)$ can be possible.
Case 2. When at least one of the subsequence still remains unbounded. Let that be the $k-$th subsequence i.e, $g(f^{n_{k}}(z_{0}))\rightarrow\infty$ i.e. $f^{n_{k}}(g(z_{0}))\rightarrow\infty$. $g(f^{m}(z_{0}))\rightarrow l_{1},$ i.e. $f^{m}(g(z_{0}))\rightarrow l_{1}$. This shows that there are at least two subsequences and in this case $g(z_{0})$ belongs to $BU(f)$. Therefore, $g(BU(f))$ is contained in either $BU(f)$ or $RP(f)$. ∎
Lemma 3.5.
$g(I(f))\subset I(f)\cup RP(f)$ for commuting functions of class $\mathcal{B}$
Proof.
Suppose $z_{0}\in I(f)$ and let $\lim_{n\rightarrow\infty}g(f^{n}(z_{0}))\rightarrow l_{0}$ (say) i.e., $f^{n}(g(z_{0}))\rightarrow l_{0}.$ This means $g(z_{0})\in K(f).$ But $z_{0}\not\in K(f)$ as it belongs to $I(f)$. So $g(z_{0})\in RP(f)$ for the above case. Now, suppose $\lim_{n\rightarrow\infty}g(f^{n}(z_{0}))\rightarrow\infty$ i.e., $f^{n}(g(z_{0}))\rightarrow\infty$. Hence, $g$ sends $I(f)$ to either $RP(f)$ or $I(f)$. So, $g(I(f))\subset I(f)\cup RP(f)$. ∎
Proposition 3.6.
$BU(f\circ g)\subset BU(f)\cup BU(g)$ for commuting functions of class $\mathcal{B}$.
Proof.
Let $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$. This implies that $z_{0}\in K(f)\cup K(g)$ or $z_{0}\in I(f)\cup I(g)$. By definition of Bungee set, there are subsequences $\{n_{k}\}$ and $\{m_{k}\}$ tending to $\infty$ satisfying
$|f^{n_{k}}(g^{n_{k}}(z_{0}))|\rightarrow\infty$
$|f^{m_{k}}(g^{m_{k}}(z_{0}))|\rightarrow l_{0}\ \ \text{for some }l_{0}.$
First, we assume that $z_{0}\in RP(g)$ (there is a clear contradiction in the $F(g)$ case because $F(g)=F(f\circ g)$), and as a result, $f^{n}(RP(g))\subset RP(g)$ and $g^{n}(f^{n}(RP(g)))$ which will be bounded. As can be seen, there are no subsequences that are unbounded. Similarly, there are no such points in $K(f)$. Therefore, $z_{0}\not\in K(f)\cup K(g)$.
Now, we assume $z_{0}\in I(g)$ which means there are two possible cases
(3.1) $|f(g^{n})(z_{0})|\rightarrow l\ \text{for some}\ l$
(3.2) $|f(g^{n})(z_{0})|\rightarrow\infty$
where $n$ denotes an arbitrary index unrelated to $n_{k}$ and $m_{k}$. If we have (3.2), then from the array of sequences argument, we see that all subsequences tend to infinity. If we have (3.1), there are further possible cases: $l\in I(f)$ or $l\in K(f)$ or $l\in BU(f)$. Again, if we have the first two cases, then we observe that entire composite sequence $(f\circ g)^{n}$ is (un)bounded. The third case is also not possible. This is because using the previous Lemma (3.4), we have $|f^{n_{k}}(g^{n_{k}})(z_{0})|$ or $|f^{m_{k}}(g^{m_{k}})(z_{0})|$ which equals $|g^{n_{k}}(f^{n_{k}})(z_{0})|$ or $|g^{m_{k}}(f^{m_{k}})(z_{0})|$ will be either completely bounded or unbounded. Therefore, $z_{0}\not\in I(g)$. On similar lines, $z_{0}\not\in I(f)$. Therefore, $z_{0}\not\in I(f)\cup I(g)$. Thus, we see that there are no points s.t, $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$. ∎
Corollary 3.7.
For commuting functions of class $\mathcal{B},(I(f)\cup I(g))\cap BU(f\circ g)=\emptyset$
Proof.
We can follow the similar lines of proof that is done in the preceeding proposition (3.6). Let $z_{0}\in I(f)$ (say). Then $g(f^{n}(z_{0}))\rightarrow\infty$ or $L$. If $g(f^{n}(z_{0}))\rightarrow\infty,$ then by array of sequences argument, we conclude that $g^{n}\circ f^{n}(z_{0})\not\rightarrow M$ for some $M$ which means that we do not have a subsequence that remains bounded upon iteration of $f\circ g$. Moreover, if $g(f^{n}(z_{0}))\rightarrow L$ then $g(z_{0})\in RP(f)$ and therefore, $f^{n}(g^{n}(f^{n}(z_{0})))\rightarrow L_{1}$ (say). Thus, we never get a subsequence that tends to $\infty$.∞ . Similarly, it can be argued when $z_{0}\in I(g)$. ∎
Revision 3.8.
Revised proof of $I(f\circ g)\subset I(f)\cup I(g)$ for commuting functions of $\mathcal{B}.$ (considering the existence of Bungee set)
Proof.
The motivation for the proof comes from [14]. Let $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ i.e., $z_{0}\in RP(f)\cup RP(g)$ or $z_{0}\in BU(f)\cup BU(g)$. By definition of escaping set: $f^{n}\circ g^{n}(z_{0})=(f\circ g)^{n}(z_{0})\rightarrow\infty$. Assume $z_{0}\in BU(f)$, then $g(BU(f))$ goes to either $BU(f)$ or $RP(f)$ (3.5). When it goes to $RP(f)$, we know $g(RP(f))$ will remain in $RP(f)$ and upon further iteration of $f\circ g$ will stay bounded i.e $(g\circ f)^{n}(z_{0})\rightarrow L$ as $n\rightarrow\infty$ which is a contradiction. If it goes to $BU(f)$ then again, its the same case as discussed before when $g$ acts on it. In the end, it will either be an entirely bounded sequence or there would be (un)bounded subsequence. The later case arises only when $g^{n}(BU(f))\subset BU(f)\ \forall n\in\mathbb{N}$ (where $n$ is some arbitrary index). But by definition escaping sets must have all the subsequences tending to infinity. Hence, we arrive at a contradiction. On similar lines, it can be shown when $z_{0}\in BU(g)$ we again arrive at a contradiction. So, $z_{0}\not\in BU(f)\cup BU(g)$.
Now the case where $z_{0}\in RP(f)\cup RP(g)$. Assuming $z_{0}\in RP(f)$ and as discussed earlier we will have all bounded subsequences of $f^{n}\circ g^{n}(z_{0})$ which is a contradiction. Similar arguments can be given when $z_{0}\in RP(g).$ Therefore $z_{0}\not\in RP(f)\cup RP(g).$ Hence, $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ is not possible. Thus, $I(f\circ g)\subset I(f)\cup I(g)$. ∎
Proposition 3.9.
$RP(f\circ g)\supset RP(f)\cup RP(g)$ for commuting functions of class $\mathcal{B}$.
Proof.
Let us consider a point $z_{0}\in RP(f)$. Then, $f^{n}(z_{0})$ is bounded and contained in $RP(f)$. Also, $g^{n}(RP(f))\subset RP(f)$, therefore $f^{n}(g^{n}(f^{n}(z_{0})))$ is also bounded as $n\rightarrow\infty.$ (Observe that, $z_{0}$ can’t be in $F(f\circ g)$ as $F(f\circ g)=F(f)$). Hence, $z_{0}\in RP(f\circ g)$ as well. On similar lines, if $w_{0}\in RP(g)$, then $w_{0}\in RP(f\circ g)$.
Now the question is can there be points not in $RP(f)\cup RP(g)$ but in $RP(f\circ g)$? To answer that let us go through (3.4) problem once again. When suppose $g(BU(f))$ goes to $RP(f)$ and upon further iteration with $(g\circ f)^{n}(z_{0})$, this will be bounded means $z_{0}\in RP(f\circ g)$. Similarly, there will be some points from $I(f)$ that will be bounded upon $g^{n}\circ f^{n}(z_{0})$ as $n\rightarrow\infty$ as discussed in (3.5). Again it has to be in $BU(g)$ and $I(g)$. So, this implies $RP(f\circ g)\supset RP(f)\cup RP(g)$. ∎
There’s a simpler proof for this which goes as follows:
$BU(f\circ g)\cup I(f\circ g)\subset BU(f)\cup BU(g)\cup I(f)\cup I(g)$
Taking complement on both sides
$BU(f\circ g)^{c}\cap I(f\circ g)^{c}\supset BU(f)^{c}\cap BU(g)^{c}\cap I(f)^{c}\cap I(g)^{c}$
(3.3) $\implies RP(f\circ g)\supset RP(f)\cup RP(g)$
Remark 3.10.
For $f\in\mathcal{B},$ we have $J(f)=\overline{I(f)}=\overline{BU(f)}=\overline{RP(f)}=I(f)\cup BU(f)\cup RP(f).$
4. Illustrations:
We now illustrate $BU(f\circ g)\subset BU(f)\cup BU(g)$ with an example.
Consider the functions $f(z)=z^{2}$ and $g=1/z^{2}$. Observe that $f\circ g(z)=1/z^{4}$. $BU(f)=\emptyset$, $BU(g)=\{z:|z|<1\cup|z|>1\}$ and $BU(f\circ g)=\{z:|z|<1\cup|z|>1\}.$ Clearly, $BU(f\circ g)\subset BU(f)\cup BU(g)$.
Now, we discuss exponential functions which are not completely exponential type. Some examples are (see [15]): let us take the case where $f=p(z)\cdot e^{g(z)}$ where $p(z),\ g(z)$ are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}$. $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
$f\circ g(z)=f(g(z))=g(z)\cdot e^{a^{(k+1)}\cdot g^{k}(z)}=aze^{a^{(k+1)}z^{k}}e^{a^{(2k+1)}z^{k}e^{ka^{(k+1)}z^{k}}}$
$g\circ f(z)=g(f(z))=a\cdot f(z)\cdot e^{a^{k}\cdot f^{k}(z)}=aze^{a^{k}z^{k}}e^{a^{2k}z^{k}e^{ka^{k}z^{k}}}.$ Now, for $f\circ g=g\circ f,$ we need to have $a^{2k+1}=a^{k}\implies a^{k}(a^{k+1}-1)=0$ which means $a$ is $(k+1)$-root of unity except ‘1’ itself because that’s trivial.
Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since a=1 is a trivial case so let us take $a=-1$, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$. Here as the orbit of any point under either of the function tends to $\infty,f(z)\rightarrow 0,g(z)\rightarrow 0$ and $z=0$ is a fixed point of $f(z),\ g(z)$ since $f(0)=0$ and $g(0)=0$. Now for the multiplier $\lambda$, $f^{\prime}(z)=e^{-z^{2}}(1-2z^{2}),f^{\prime}(0)=1$ which is an indifferent fixed point. And $11^{1}=$ 1 means $0\in J(f)$ or $J(g)$ (since both of them are equal). As $\infty$ is a special case and belongs to $J(f)$ so, this proves that a sequence which tends to $\infty$ if it asymptotes to a finite asymptotic value which is a member of the Julia set of $f$. Moreover, the function which are of the form $f=p(z)e^{g(z)}$ which asymptotes because it has exponential term, means the asymptotic value is ${}^{\prime}0^{\prime}$ and it is in $J(f)$ (rationally-indifferent fixed point).
Now some illustrations for the Bungee set:
Let $f(z)=ze^{z^{2}};\ g(z)=-ze^{z^{2}}$ $\implies f^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}\ \&\ g^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}.$ $f\circ g=-ze^{z^{2}}e^{z^{2}e^{2z^{2}}}=g\circ f$. And $(f\circ g)^{2}=g^{4}=f^{4}$(Because as $(f\circ g)^{2}$ is done it does away with the negative term and rest terms are exactly same). So, $BU(f^{4})=BU(g^{4})=BU((f\circ g)^{2})\implies BU(f)=BU(g)=BU(f\circ g)$ (NOTE: If we have $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$, they also commute). We may also take $f(z)=ze^{a^{(k+1)}z^{k}}$, $g(z)=aze^{a^{(k+1)}z^{k}}$ and $a^{(k+1)}=1$ for that matter). And this complies with the union relation with the composition.
Now, for the general case: Since $a$ is $k$-th root of unity, means $a^{k}=1$. So, $f(z)=ze^{a^{k}z^{k}}=ze^{z^{k}}$ and $g(z)=aze^{z^{k}}$. Now $g^{k}(z)=a^{k}\cdot f^{k}(z)=f^{k}(z)\implies BU(f)=BU(g)$. Also, $(f\circ g)^{k}(z)=a^{k}f^{2k}(z)=f^{2k}(z)$. So, $BU(f\circ g)=BU(f)=BU(g)$. Analogously, if we take $f(z)=ze^{a^{(k+1)}z^{k}}$ and $\ g(z)=af(z)=aze^{a^{(k+1)}z^{k}}$, where again $a^{k+1}=1$, we can simplify this further as $f(z)=ze^{az^{k}}$ and $g(z)=aze^{az^{k}}$, now $g^{2}(z)=af(af(z)))=a^{2}f^{2}(z)$. So, like this $g^{k+1}(z)=a^{(k+1)}f^{k+1}(z)$. Also $f\circ g(z)=af^{2}(z)=g\circ f(z)$. Now, if we do $(f\circ g)^{2}(z)=a^{2}f^{4}(z)$. So, $(f\circ g)^{k+1}(z)=a^{(k+1)}f^{2(k+1)}(z)=f^{2(k+1)}(z)$. Here also we can see that $BU(f\circ g)=BU(f)=BU(g)$.
Thus, functions which are of the form $f(z)=ze^{a^{k}z^{k}}$ where $a$ is $k$-th root of unity or $f(z)=ze^{a^{k+1}z^{k}}$, where $a$ is $(k+1)$-th root of unity and $g=af$ have the property that their bungee sets are same. Moreover, their escaping set and repelling periodic point sets are also the same. $J(f)=J(g)$ also holds true because we are taking $f=p(z)e^{g(z)}$ as discussed in [15].
5. Conclusion:
We have shown that entire functions with no finite asymptotic values which are also commuting i.e, $f\circ g=g\circ f$ satisfies:
$K(f)=K(g),\ I(f)=I(g),\ BU(f)=BU(g)$ and therefore $K(f\circ g)=K(f),\ BU(f\circ g)=BU(f),\ I(f\circ g)=I(f)$.
And for the functions of class $\mathcal{B}$ which are commuting
$RP(f\circ g)\supset RP(f)\cup RP(g)$;
$BU(f\circ g)\subset BU(f)\cup BU(g)$;
$I(f\circ g)\subset I(f)\cup I(g)$.
There are some other special cases where we have equality among them like when $f$ is an exponential family and cases like $g=f+C$ where $C\in\mathbb{C}$ which are periodic translate of $f$. Also, when functions are from class $\mathcal{B}$, $(BU(f)\cup BU(g))\cap RP(f\circ g)\neq\emptyset$ because $g(BU(f))\subset RP(f)\cup BU(f)$. Similarly, $(I(f)\cup I(g))\cap RP(f\circ g)\neq\emptyset$ because $g(I(f))\subset RP(f)\cup I(f)$. Also, there are no points which are in $I(f)\cup I(g)$ as well as in $BU(f\circ g)$.